# An Objective Improvement Approach

### v2

@@ 1. Introduction
Various classes of graph games are characterised by the objective of the players, for instance in parity games the objective is to optimise the parity of the dominating colour occurring infinitely often, while in discounted and mean-payoff games the objective of the players is to minimise resp. maximise the discounted and limit-average sum of the colours.
Solving graph games is the central and most expensive step in many model checking [Koz83, EJS93, Wil01, dAHM01, AHK02, SF06a], satisfiability checking [Koz83, Var98, Wil01], and synthesis [Pit06, SF06b] algorithms. Progress in algorithms for solving graph games will therefore allow for the development of more efficient model checkers and contribute to bringing synthesis techniques to practice.
There is a hierarchy among the graph games mentioned earlier, with simple and well known reductions from parity games to mean payoff games, from mean-payoff games to discounted payoff games, and from discounted payoff games to simple stochastic games like the ones from [ZP96], while no reductions are known in the other direction. Therefore, one can solve instances of all these games by using an algorithm for stochastic games. All of these games are in UP and co-UP [Jur98], while no tractable algorithm is known.
@@
Interestingly, the one class of efficient techniques for solving parity games that does not (yet) have a quasi-polynomial approach is strategy improvement algorithms [Lud95, Pur95, VJ00, BV07, Sch08, Fea10, STV15], a class of algorithms closely related to the Simplex for linear programming, known to perform well in practice. Most of these algorithms reduce to mean [BV07, Sch08, STV15, BDM20] or discounted [Lud95, Pur95, FGO20, Koz21] payoff games.
With the exception of the case in which the fixed-point of discounted payoff games is explicitly computed [ZP96], all these algorithms share a disappointing feature: they are inherently non-symmetric approaches for solving an inherently symmetric problem. However, some of these approaches have a degree of symmetry. Recursive approaches treat even and odd colours symmetrically, one at a time, but they treat the two players very differently for a given colour. Symmetric strategy improvement [STV15] runs a strategy improvement algorithms for both players in parallel, using the intermediate results of each of them to inform the updates of the other, but at heart, these are still two intertwined strategy improvement algorithms that, individually, are not symmetric. This is in due to the fact that applying strategy improvement itself symmetrically can lead to cycles [Con93].
The key contribution of this paper is to devise a new class of algorithms to solve discounted payoff games, which is entirely symmetric. Like strategy improvement algorithms, it seeks to find co-optimal strategies, and improves strategies while they are not optimal. In order to do so, however, it does not distinguish between the strategies of the two players. This seems surprising, as maximising and minimising appear to pull in opposing directions.
Similar to strategy improvement approaches, the new objective improvement approach turns the edges of a game into constraints (here called inequations), and minimises an objective function. However, while strategy improvement algorithms take only the edges in the strategy of one player (and all edges of the other player) into account and then finds the optimal response by solving the resulting one player game, objective improvement always takes all edges into account. The strategies under consideration then form a subset of the inequations, and the goal would be to make them sharp (i.e. strict, tight, satisfied as equations), which only works when both strategies are optimal. When they are not, then there is some offset for each of the inequations, and the objective is to reduce this offset in every improvement step. This treats the strategies of both players completely symmetrically.
@@
The paper is organised as follows. After the preliminaries in Section 2, we start by outlining our method and use a simple game to explain it in Section 3. We then formally introduce our objective improvement algorithm in Section 4, keeping the question of how to choose better strategies abstract. Section 5 discusses how to find better strategies, and Section 6 provides an experimental evaluation of the algorithm. We finally wrap up with a discussion of our results in Section 7.
@@ 2. Preliminaries
A discounted payoff game (DPG) is a tuple $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$, where $V=V_{\min}\cup V_{\max}$ are the vertices of the game, partitioned into two disjoint sets $V_{\min}$ and $V_{\max}$, such that the pair $(V,E)$ is a finite directed graph without sinks. The vertices in $V_{\max}$ (resp, $V_{\min}$) are controlled by Player Max or maximiser (resp, Player Min or minimiser) and $E\subseteq V\times V$ is the edge relation. Every edge has a weight represented by the function $w:E\to\mathbb{R}$, and a discount factor represented by the function $\lambda:E\to[0,1)$. When the discount factor is uniform, i.e. the same for every edge, it is represented by a constant value $\lambda\in[0,1)$. For ease of notation, we write $w_{e}$ and $\lambda_{e}$ instead of $w(e)$ and $\lambda(e)$. A play on $\mathcal{G}$ from a vertex $v$ is an infinite path, which can be represented as a sequence of edges $\rho=e_{0}e_{1}e_{2}\ldots$ such that, for every $i\in\mathbb{N}$, $e_{i}=(v_{i},v_{i+1})\in E$ and $v_{0}=v$. By $\rho_{i}$ we refer to the $i$-th edge of the play. The outcome of a discounted game $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ for a play $\rho$ is $\mathsf{out}(\rho)=\sum_{i=0}^{\infty}w_{e_{i}}\prod_{j=0}^{i-1}\lambda_{e_{j}}$. For games with a constant discount factor, this simplifies into $\mathsf{out}(\rho)=\sum_{i=0}^{\infty}w_{e_{i}}\lambda^{i}$.
A positional strategy for Max is a function $\sigma_{\max}:V_{\max}\to V$ that maps each Max vertex to a vertex according to the set of edges, i.e. $(v,\sigma_{\max}(v))\in E$. Positional Min strategies are defined accordingly, and we call the set of positional Min and Max strategies $\Sigma_{\min}$ and $\Sigma_{\max}$, respectively.
@@
A positional maximiser (resp. minimiser) strategy $\sigma$ is said to be optimal if, and only if, $\mathsf{val}_{\mathcal{G}}(v)=w_{(v,\sigma(v))}+\lambda_{(v,\sigma(v))}\mathsf{val}_{\mathcal{G}}(\sigma(v))$ holds for all maximiser (resp. minimiser) vertices.
@@
If both of these strategies are optimal, we call $\sigma$ a joint co-optimal strategy. This is the case if, and only if, $\mathsf{val}_{\mathcal{G}}=\mathsf{val}_{\sigma}$ holds.
@@
For co-optimal strategies $\sigma$, $\mathsf{val}_{\sigma}$ that is equal to $\mathsf{val}_{\mathcal{G}}$ is clearly a solution, as all of these inequations are satisfied, while for a pair of strategies $\sigma$ that is not co-optimal, $\mathsf{val}_{\sigma}$ is not a solution.
@@ 2.1. Simplex method
A common way to compute $\mathsf{val}_{\mathcal{G}}$ is to create a linear programming instance from $\mathcal{G}$. In this paragraph we provide the background for solving a linear programming instance and outline the Simplex method that is the standard algorithm used to find a solution.
@@
A linear programming problem requires to find a (solution) vector $x\in\mathbb{R}^{n}$ that maximises (or minimises) $c\cdot x$, where $c\in\mathbb{R}^{n}$ is a vector of constants, such that $A\cdot x\leq b$ (or $\geq b$), where $A\in\mathbb{R}^{m\times n}$ is a coefficient matrix and $b\in\mathbb{R}^{m}$ is a vector of constants. Every linear programming instance can be expressed in a simpler standard form by applying transformations to the objective function and to the constraints. A standard form requires to find $x$ that minimises $c\cdot x$ such that $A\cdot x=b$ and $x\geq 0$, with $b$ positive. Given a game $\mathcal{G}$, we can define an objective function, the input matrix $A$, and the vectors $b$ and $c$, such that $x=\mathsf{val}_{\mathcal{G}}$, in the following way. $A$ is composed of $m$ rows and $n$ columns where $m$ is the number of edges and $n$ the number of vertices in $\mathcal{G}$. Since the constrains in $A$ represent the edges, that are binary relations, only two columns in each row contain a non-zero value (with the exception of self-loops that have only one non-zero coefficient). The vector $b$ contains the weight of the edges. Hence, every row $i$ can be expressed as $a_{ij}+a_{ik}\geq b_{i}$, if $e=(j,k)$ is an edge of $\mathcal{G}$, $j$ is a maximiser vertex (otherwise the constraint is $\leq b_{i}$), and $b_{i}=w_{e}$. Finally, $c$ is the unit vector with a positive sign for the maximiser vertices and negative for the minimiser, and the objective function maximises the sum of the values of $x$.
@@
From a graphical point of view, the set of costraints, i.e. the inequations represented by $A$ and $b$, define a region called polytope (or: simplex; hence the name of the approach) in the space $\mathbb{R}^{n}$. Any point that lies within the polytope is a solution. The objective function, however, provides an additional vector that shows the direction of better solutions. As a result, if there is an optimal solution (and there always is an optimal solution in the linear programming problems we consider in this paper), then there is an optimal solution that lies on a corner of the polytope. This observation allows to restrict the search space for optimal solutions to the corners of the polytope. In fact, it suffices to move along the edges of the polytope to find and optimal solution.
From an algebraic point of view, each vector $x$ that corresponds to a corner of the polytope satisfies $n$ constraints (vectors) linearly independent of the linear programming problem. Such $x$ is called basis since every constraint of the problem can be written as a unique combination of the $n$ constraints satisfied by $x$. The associated set of constraints to the basis is called $B$. Hence, to find the optimal solution of the linear programming problem one has to find the basis (admissible solution) that minimises the objective function (optimality).
@@
The Simplex method assumes that the problem is in the standard form and solves it by finding a basis and then by moving to an adjacent basis that improves the value of the objective function. Two basis are adjacent if in the graphical representation they correspond to adjacent corners. This implies that the set of constraints they satisfy differ for one single vector. The algorithm stops when the adjacent bases do not improve the value of the objective function of the basis currently selected.
@@
1 $B\leftarrow{}$ random initial basis matrix
@@
    foreach $j$ non basis index do
       3 $c^{\prime}_{j}\leftarrow{}$ reduced cost over objective function
@@
   6 $B\leftarrow{}\mathsf{Pivot}(A,b,c,j)$
@@ 7 $x\leftarrow{}\mathsf{Update}(B,c)$
The procedure is reported in Algorithm 1. At Line 1, an initialisation step selects a first basis. The “for each” loop at Line 3 computes the difference of the current objective function value and the one obtained by selecting a different constraint $j$. In case all these differences are positive, then the current solution cannot be improved and it is returned at Line 4. Otherwise, one of the improving adjacent corner is selected (Line 5). The move from one basis to an adjacent one is called pivoting operation (Line 6). Once the basis matrix has been changed, the solution vertex can be updated (Line 7).
@@ 3. Outline and Motivation Example
We start by considering the simple discounted payoff game of Figure 1, assuming that it has some uniform discount factor $\lambda\in[0,1)$. In this game, the minimiser (who owns the right vertex, marked by a square), has only one option: she always has to use the self-loop, which earns her an immediate reward of $1$. The overall reward the minimiser reaps for a play that starts in her vertex is therefore $1+\lambda+\lambda^{2}+\ldots=\frac{1}{1-\lambda}$.
@@
The optimal decision of the maximiser is therefore to move on the first time, which yields the maximal reward of $\frac{\lambda}{1-\lambda}$. Every vertex $v$ has some outgoing edge(s) $e=(v,v^{\prime})$ where $\mathsf{val}(v)=w_{e}+\lambda_{e}\mathsf{val}(v^{\prime})$ holds [ZP96]; these edges correspond to the optimal decisions for the respective player.
For our running example game of Figure 1 with a fixed discount factor $\lambda\in[0,1)$, the constraints a solution must satisfy are:
@@
The unique valuation that satisfies these inequations and produces a sharp inequation (i.e. satisfied as equation) for some outgoing edge of each vertex assigns $\mathsf{val}(a)=\frac{1}{1-\lambda}$ and $\mathsf{val}(b)=\frac{\lambda}{1-\lambda}$. This valuation also defines the optimal strategies of the players (to stay for the minimiser, and to move on for the maximiser).
@@
We discuss a symmetric approach to find this unique valuation. Our approach adjusts linear programming in a natural way that treats both players symmetrically: we maintain the set of inequations for the complete time, while approximating the goal of “one equality per vertex” by the objective function. To do that, we initially fix an arbitrary outgoing edge for every vertex (a strategy), and minimise the sum of the distances between the left and right side of the inequations defined by these edges, which we call the offset of this edge. This means, for an edge $e=(v,v^{\prime})$, to minimise the difference between $\mathsf{val}(v)$ (left side of the inequation) and $w_{e}+\lambda_{e}\mathsf{val}(v^{\prime})$ (right side).
To make this clear, we consider again the example of Figure 1 and use both self loops as the strategies for the players fixed at the beginning in our running example. The offset for the selected outgoing edge of the minimiser vertex $a$ is equal to $1-(1-\lambda)\mathsf{val}(a)$, while the offset for the selected outgoing edge of the maximiser vertex $b$ is equal to $(1-\lambda)\mathsf{val}(b)$. The resulting overall objective consists, therefore, in minimising the value $1-(1-\lambda)\mathsf{val}(a)+(1-\lambda)\mathsf{val}(b)$.
This term is always non-negative, since it corresponds to the sum of the edges contributions that are all non-negative. Moreover, when only optimal strategies are selected to form this objective function, the value 00 can be taken, and where it is taken, it defines the correct valuation.
@@
Thus, in this toy example, while selecting the wrong edge cannot result in the objective function taking the value 00, we still found the optimal solution. In general, we might need to update the objective function. To update the objective function, we change the outgoing edges of some (or all) vertices, such that the overall value of the objective function goes down. Note that this can be done not only when the linear program returns an optimal solution, but also during its computation. For example, when using a simplex method, updating the objective function can be used as an alternative pivoting rule at any point of the traversal of the simplex.
In general, the valuation returned as solution is computed by using objective functions based on sub-optimal strategies. The simplest way of seeing non-optimal strategies is to use different discount factors [^2] Note that we can also replace the transitions with a smaller discount factor by multiple transitions with a larger discount factor. This would allow for keeping the discount factor uniform, but needlessly complicate the discussion and inflate the size of the example. for the game of Figure 1, let’s say $\frac{1}{3}$ for the self-loop of the maximiser vertex and $\frac{2}{3}$ for the other two transitions, so that the three equations are: (1) $\mathsf{val}(a)\leq 1+\frac{2}{3}\mathsf{val}(a)$, (2) $\mathsf{val}(b)\geq\frac{2}{3}\mathsf{val}(b)$, and (3) $\mathsf{val}(b)\geq\frac{1}{3}\mathsf{val}(a)$. Making the adjusted inequations (2) and (3) sharp still results in the values $\mathsf{val}(a)=\mathsf{val}(b)=0$, and the objective function still takes the value of $1$. While making inequations (1) and (3) sharp provides the values $\mathsf{val}(a)=3$ and $\mathsf{val}(b)=2$; the objective function takes the value $\frac{4}{3}$ at this point. Finally, if we consider the last combination, making (1) and (2) sharp still conflicts with inequation (3).

### v3

@@ 1. Introduction
Various classes of graph games are characterised by the objective of the players, for instance, in parity games the objective is to optimise the parity of the dominating colour occurring infinitely often, while in discounted and mean-payoff games the objective of the players is to minimise resp. maximise the discounted and limit-average sum of the colours.
Solving graph games is the central and most expensive step in many model checking [Koz83, EJS93, Wil01, dAHM01, AHK02, SF06a], satisfiability checking [Koz83, Var98, Wil01], and synthesis [Pit06, SF06b] algorithms. Progress in algorithms for solving graph games will, therefore, allow for the development of more efficient model checkers and contribute to bringing synthesis techniques to practice.
There is a hierarchy among the graph games mentioned earlier, with simple and well-known reductions from parity games to mean payoff games, from mean-payoff games to discounted payoff games, and from discounted payoff games to simple stochastic games such as the ones from [ZP96], while no reductions are known in the other direction. Therefore, one can solve instances of all these games by using an algorithm for stochastic games. All of these games are in UP and co-UP [Jur98], while no tractable algorithm is known.
@@
Interestingly, the only class of efficient techniques for solving parity games that does not (yet) have a quasi-polynomial approach is strategy improvement algorithms [Lud95, Pur95, VJ00, BV07, Sch08, Fea10, STV15], a class of algorithms closely related to the Simplex for linear programming, known to perform well in practice. Most of these algorithms reduce to mean [BV07, Sch08, STV15, BDM20] or discounted [Lud95, Pur95, FGO20, Koz21] payoff games.
With the exception of the case in which the fixed-point of discounted payoff games is explicitly computed [ZP96], all these algorithms share a disappointing feature: they are inherently non-symmetric approaches for solving an inherently symmetric problem. However, some of these approaches have a degree of symmetry. Recursive approaches treat even and odd colours symmetrically, one at a time, but they treat the two players very differently for a given colour. Symmetric strategy improvement [STV15] runs a strategy improvement algorithms for both players in parallel, using the intermediate results of each of them to inform the updates of the other, but at heart, these are still two intertwined strategy improvement algorithms that, individually, are not symmetric. This is due to the fact that applying strategy improvement itself symmetrically can lead to cycles [Con93].
The key contribution of this paper is to devise a new class of algorithms to solve discounted payoff games, which is entirely symmetric. Like strategy improvement algorithms, it seeks to find co-optimal strategies and improves strategies while they are not optimal. However, in order to do so, it does not distinguish between the strategies of the two players. This seems surprising, as maximising and minimising appear to pull in opposing directions.
Similarly to strategy improvement approaches, the new objective improvement approach turns the edges of a game into constraints (here called inequations) and minimises an objective function. However, while strategy improvement algorithms take only the edges in the strategy of one player (and all edges of the other player) into account and then finds the optimal response by solving the resulting one-player game, objective improvement always takes all edges into account. The strategies under consideration then form a subset of the inequations, and the goal would be to make them sharp (i.e., strict, tight, satisfied as equations), which only works when both strategies are optimal. When they are not, then there is some offset for each of the inequations, and the objective is to reduce this offset in every improvement step. This treats the strategies of both players completely symmetrically.
@@
The paper is organised as follows. After the preliminaries in Section 2, we start by outlining our method and using a simple game to explain it in Section 3. We then formally introduce our objective improvement algorithm in Section 4, keeping the question of how to choose better strategies abstract. Section 5 discusses how to find better strategies and Section 6 provides an experimental evaluation of the algorithm. We finally wrap up with a discussion of our results in Section 7.
@@ 2. Preliminaries
A discounted payoff game (DPG) is a tuple $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$, where $V=V_{\min}\cup V_{\max}$ are the vertices of the game, partitioned into two disjoint sets $V_{\min}$ and $V_{\max}$, such that the pair $(V,E)$ is a finite directed graph without sinks. The vertices in $V_{\max}$ (resp, $V_{\min}$) are controlled by Player Max or maximiser (resp, Player Min or minimiser) and $E\subseteq V\times V$ is the edge relation. Every edge has a weight represented by the function $w:E\to\mathbb{R}$, and a discount factor represented by the function $\lambda:E\to[0,1)$. When the discount factor is uniform, i.e., the same for every edge, it is represented by a constant value $\lambda\in[0,1)$. For ease of notation, we write $w_{e}$ and $\lambda_{e}$ instead of $w(e)$ and $\lambda(e)$. A play on $\mathcal{G}$ from a vertex $v$ is an infinite path, which can be represented as a sequence of edges $\rho=e_{0}e_{1}e_{2}\ldots$ such that, for every $i\in\mathbb{N}$, $e_{i}=(v_{i},v_{i+1})\in E$ and $v_{0}=v$. By $\rho_{i}$ we refer to the $i$-th edge of the play. The outcome of a discounted game $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ for a play $\rho$ is $\mathsf{out}(\rho)=\sum_{i=0}^{\infty}w_{e_{i}}\prod_{j=0}^{i-1}\lambda_{e_{j}}$. For games with a constant discount factor, this simplifies into $\mathsf{out}(\rho)=\sum_{i=0}^{\infty}w_{e_{i}}\lambda^{i}$.
A positional strategy for Max is a function $\sigma_{\max}:V_{\max}\to V$ that maps each Max vertex to a vertex according to the set of edges, i.e., $(v,\sigma_{\max}(v))\in E$. Positional Min strategies are defined accordingly, and we call the set of positional Min and Max strategies $\Sigma_{\min}$ and $\Sigma_{\max}$, respectively.
@@
A positional maximiser (resp. minimiser) strategy $\sigma$ is said to be optimal if and only if $\mathsf{val}_{\mathcal{G}}(v)=w_{(v,\sigma(v))}+\lambda_{(v,\sigma(v))}\mathsf{val}_{\mathcal{G}}(\sigma(v))$ holds for all maximiser (resp. minimiser) vertices.
@@
If both of these strategies are optimal, we call $\sigma$ a joint co-optimal strategy. This is the case if, and only if, $\mathsf{val}_{\mathcal{G}}=\mathsf{val}_{\sigma}$ holds [^1] If $\mathsf{val}_{\sigma}=\mathsf{val}_{\mathcal{G}}$ holds, then, for every vertex $v$ and for $v^{\prime}=\sigma(v)$, $\mathsf{val}_{\sigma}(v)=w_{(v,v^{\prime})}+\lambda_{(v,v^{\prime})}\mathsf{val}_{\sigma}(v^{\prime})$ holds. At the same time, for all maximiser vertices $v$ and edges $(v,v^{\prime\prime})$, it then holds that $\mathsf{val}_{\sigma}(v)\leq w_{(v,v^{\prime\prime})}+\lambda_{(v,v^{\prime\prime})}\mathsf{val}_{\sigma}(v^{\prime\prime})$, while, for all minimiser vertices $v$ and edges $(v,v^{\prime\prime})$, $\mathsf{val}_{\sigma}(v)\geq w_{(v,v^{\prime\prime})}+\lambda_{(v,v^{\prime\prime})}\mathsf{val}_{\sigma}(v^{\prime\prime})$ holds. This implies that the minimiser and maximiser strategies constituting $\sigma$ are co-optimal..
@@
For co-optimal strategies $\sigma$, $\mathsf{val}_{\sigma}$ that is equal to $\mathsf{val}_{\mathcal{G}}$ is clearly a solution, as all of these inequations are satisfied, while for a pair of strategies $\sigma$ that is not co-optimal, $\mathsf{val}_{\sigma}$ is not a solution [^3] If $\sigma$ is not co-optimal, then there is a maximiser vertex $v$ and an edge $e=(v,v^{\prime})$ such that $\mathsf{val}(v)<w_{e}+\lambda_{e}\mathsf{val}(v^{\prime})$ or there is a minimiser vertex $v$ and an edge $e=(v,v^{\prime})$ such that $\mathsf{val}(v)>w_{e}+\lambda_{e}\mathsf{val}(v^{\prime})$. In both cases, the inequation for this particular edge $e$ is violated..
@@ 2.1. Simplex method
A common way to compute $\mathsf{val}_{\mathcal{G}}$ is to create a linear programming instance from $\mathcal{G}$. In this paragraph, we provide the background for solving a linear programming instance and outline the Simplex method that is the standard algorithm used to find a solution.
@@
A linear programming problem requires to find a (solution) vector $x\in\mathbb{R}^{n}$ that maximises (or minimises) $c\cdot x$, where $c\in\mathbb{R}^{n}$ is a vector of constants, such that $A\cdot x\leq b$ (or $\geq b$), where $A\in\mathbb{R}^{m\times n}$ is a coefficient matrix and $b\in\mathbb{R}^{m}$ is a vector of constants. Given a game $\mathcal{G}$, we can define an objective function represented by the vector c, the input matrix $A$, and the vector $b$, such that the solution $x$ to the linear programming instance corresponds to $\mathsf{val}_{\mathcal{G}}$, as follows. In this translation from $\mathcal{G}$ to the linear programming instance, the matrix $A$ is composed of $m$ rows and $n$ columns where $m$ is the number of edges and $n$ the number of vertices in $\mathcal{G}$. Since the constrains in $A$ represent the edges, that are binary relations, only two columns in each row contain a nonzero value (with the exception of self-loops that have only one nonzero coefficient). The vector $b$ contains the weight of the edges. Hence, every row $i$ can be expressed as $a_{ij}\cdot x_{j}+a_{ik}\cdot x_{k}\geq b_{i}$, if $e=(j,k)$ is an edge of $\mathcal{G}$ and $j$ is a maximiser vertex (otherwise the constraint is $\leq b_{i}$), with $a_{ij}=1$, $a_{ik}=\lambda_{e}$, and $b_{i}=w_{e}$. Finally, $c$ is the unit vector with a positive sign for the maximiser vertices and negative sign for the minimiser, and the objective function maximises the sum of the values of $x$. Every linear programming instance can be expressed in a simpler standard form by applying transformations to the objective function and to the constraints. A standard form requires to find $x$ that minimises $c\cdot x$ such that $B\cdot x=b$ and $x\geq 0$, with $b$ positive and $B$ a sub-matrix of $n$ linear independent rows vectors of $A$. Such a form can always be obtained by introducing new variables, called slack, that transform the inequations into equations. Note that none of the algorithms for solving DPGs requires standard form instances (as described in the next section, the strategy improvement approach solves single-player DPGs with a fixed objective function, while the approach we present in this work solves two-players DPGs but on a different objective function at each iteration), however, to describe the Simplex method, in this section we assume the linear programming instance is translated in standard form.
@@
From a graphical point of view, the set of costraints, i.e., the rows of $B$ and $b$, define a region called polytope (or: simplex; hence the name of the approach) in the space $\mathbb{R}^{n}$. Any point that lies within the polytope is a solution. The objective function, however, provides an additional vector that shows the direction of better solutions. As a result, if there is an optimal solution (and there always is an optimal solution in the linear programming problems we consider in this paper), then there is an optimal solution that lies on a corner of the polytope. This observation allows to restrict the search space for optimal solutions to the corners of the polytope. In fact, it suffices to move along the edges of the polytope to find an optimal solution.
From an algebraic point of view, each non-zero variable in a solution $x$ is active. The subset of $n$ constraints of $A$ selected by $B$ is called basis, since every constraint of $A$ can be written as a unique combination of the $n$ constraints in $B$. If two bases $B$ or solutions $x$ differ for one active variable, then they are adjacent.
@@
The Simplex method solves instances in standard form by finding a solution and then by moving to an adjacent solution that improves the value of the objective function. In the graphical representation, two adjacent solutions lie on adjacent corners. The algorithm stops when the adjacent solutions do not improve the value of the objective function of the one currently selected.
@@
1 $B\leftarrow{}$ random initial basis matrix from $A$
@@
    foreach $j$ non-basis index do
       3 $c^{\prime}_{j}\leftarrow{}$ reduced cost over objective function for index j
@@
   6 Select $k$ such that $k$ is the min ratio test index
   7 $B\leftarrow{}\mathsf{Pivot}(A,b,c,j,k)$
@@ 8 $x\leftarrow{}\mathsf{Update}(B,c)$
The procedure is reported in Algorithm 1. At Line 1, an initialisation step selects a first basis (a subset of $n$ rows from $A$ which are linearly independent, thus $B$ is a square $n$-matrix). The “for each” loop at Line 3 computes, for every index $j$ of a variable that does not belong (i.e., not active) to the current basis $B$, the difference between the current objective function value and the one obtained by selecting a different variable $j$. Such a difference is called reduced cost. In case all these differences are non-negative, that is, when the vector $c^{\prime}$ composed of the $c^{\prime}_{j}$ costs is non-negative, then the current solution cannot be improved and it is returned at Line 4. The index $j$ corresponds to a column of $B$ and identifies the candidate variable to enter the basis. Otherwise, at Line 5, one variable from the current basis is selected to be replaced by $j$. The leaving variable $k$ is found by computing for each constraint (row of $B$) the ratio between the constant value from $b$ corresponding to the current row and the coefficient of the candidate variable to join the basis. If the $k$-th row has such a minimum value, then one of the improving adjacent corners is selected as solution by replacing $k$ with $j$. This check is called minimum ratio test, while the move from a basis to an adjacent one is called pivoting operation (Line 7) and consists in column operations over $B$ that maintain the linear independence of the matrix. Once the basis matrix has been changed, the solution vertex can be updated (Line 8).
@@ 3. Outline and Motivation Example
We start by considering the simple discounted payoff game of Figure 1, assuming that it has some uniform discount factor $\lambda\in[0,1)$. In this game, the minimiser (who owns the right vertex, depicted as a square) has only one option: she always has to use the self-loop, which earns her an immediate reward of $1$. The overall reward the minimiser reaps for a play that starts in her vertex is therefore $1+\lambda+\lambda^{2}+\ldots=\frac{1}{1-\lambda}$.
@@
The optimal decision of the maximiser is, therefore, to move on the first time, which yields the maximal reward of $\frac{\lambda}{1-\lambda}$. Every vertex $v$ has some outgoing edge(s) $e=(v,v^{\prime})$ where $\mathsf{val}(v)=w_{e}+\lambda_{e}\mathsf{val}(v^{\prime})$ holds [ZP96]; these edges correspond to the optimal decisions for the respective player.
For our running example game of Figure 1 with a fixed discount factor $\lambda\in[0,1)$, the constraints that a solution must satisfy are:
@@
The unique valuation that satisfies these inequations and produces a sharp inequation (i.e., satisfied as equation) for some outgoing edge of each vertex assigns $\mathsf{val}(a)=\frac{1}{1-\lambda}$ and $\mathsf{val}(b)=\frac{\lambda}{1-\lambda}$. This valuation also defines the optimal strategies of the players (to stay for the minimiser, and to move on for the maximiser).
@@
We discuss a symmetric approach to find this unique valuation. Our approach adjusts linear programming in a natural way that treats both players symmetrically: we maintain the set of inequations for the complete time, while approximating the goal of “one equality per vertex” by the objective function. To do that, we initially fix an arbitrary outgoing edge for every vertex (a strategy), and minimise the sum of the distances between the left and right sides of the inequations defined by these edges, which we call the offset of this edge. This means, for an edge $e=(v,v^{\prime})$, to minimise the difference between $\mathsf{val}(v)$ (left side of the inequation) and $w_{e}+\lambda_{e}\mathsf{val}(v^{\prime})$ (right side).
To make this clear, we consider again the example of Figure 1 and use both self-loops as the strategies for the players fixed at the beginning in our running example. The offset for the selected outgoing edge of the minimiser vertex $a$ is equal to $1-(1-\lambda)\mathsf{val}(a)$, while the offset for the selected outgoing edge of the maximiser vertex $b$ is equal to $(1-\lambda)\mathsf{val}(b)$. The resulting overall objective consists, therefore, in minimising the value $1-(1-\lambda)\mathsf{val}(a)+(1-\lambda)\mathsf{val}(b)$.
This term is always non-negative, since it corresponds to the sum of the edges’ contributions that are all non-negative. Moreover, when only optimal strategies are selected to form this objective function, the value 00 can be taken, and where it is taken, it defines the correct valuation.
@@
Thus, in this toy example, while selecting the wrong edge cannot result in the objective function taking the value 00, we still found the optimal solution. In general, we might need to update the objective function. To update the objective function, we change the outgoing edges of some (or all) vertices so that the overall value of the objective function goes down. Note that this can be done not only when the linear program returns an optimal solution but also during its computation. For example, when using a simplex method, updating the objective function can be used as an alternative pivoting rule at any point during the traversal of the simplex.
In general, the valuation returned as solution is computed by using objective functions based on strategies that are not necessarily optimal. We now adjust the example so that choosing both self-loops is not only non-optimal, but also does not define the correct valuation. To this end, we use different discount factors [^4] Note that we can also replace the transitions with a smaller discount factor by multiple transitions with a larger discount factor. This would allow for keeping the discount factor uniform, but needlessly complicate the discussion and inflate the size of the example. for the game of Figure 1: we choose $\frac{1}{3}$ for the self-loop of the maximiser vertex and $\frac{2}{3}$ for the other two transitions, so that the three resulting inequations are (1) $\mathsf{val}(a)\leq 1+\frac{2}{3}\mathsf{val}(a)$; (2) $\mathsf{val}(b)\geq\frac{1}{3}\mathsf{val}(b)$; and (3) $\mathsf{val}(b)\geq\frac{2}{3}\mathsf{val}(a)$. Making the adjusted inequations (2) and (3) sharp still results in the values $\mathsf{val}(a)=\mathsf{val}(b)=0$, and the objective function still takes the value of $1$. While making inequations (1) and (3) sharp provides the values $\mathsf{val}(a)=3$ and $\mathsf{val}(b)=2$; the objective function takes the value $\frac{4}{3}$ at this point. Finally, if we consider the last combination, making (1) and (2) sharp still conflicts with inequation (3).